The company wishes to increase its total dollar contribution margin by 65% in 2021. Determine by how much it will need to increase its sales if all other factors remain constant.Montreal Seating Co., a manufacturer of chairs, had the following data for 2020: Sales 2,930 units Sales price $40 per unit V
Answer #1

Contribution margin = Sales price – Variable cost = 40 – 28 = 12

Total contribution margin = Contribution margin × Sales

                                                = 12 × 2,930

                                                = 35,160

Now, this is to be increased by 65% or 0.65.

Increased total contribution margin = Total contribution margin × (1 + increased rate)

                                                            = 35,160 × (1 + 0.65)

                                                            = 35,160 × 1.65

                                                            = 58,014

Let the increased sales is Q.

Now, by the use of total contribution margin formula,

Increased total contribution margin = Contribution margin × Sales

58,014 = 12 × Q

58,014 / 12 = Q

Q = 4,834.50

Need to increase sales by (4,834.50 – 2,930 =) 1,904.50 units (Answer)

Note: fixed cost is not relevant here.

> Or: Contribution margin in dollars = (sales x sales price) – (sales x variable cost)
Contribution margin in dollars = ($2,930 x $40) – ($2,930 x $28)
= $117,200 – $82,040 = $35,160
Now, this is to be increased by 65% or 0.65 = 1+0.65 = 1.65
$35,160 x 1,65 = 58,014
Let the increased sales is Q.
Now, by the use of total contribution margin formula,
Increased total contribution margin = Contribution margin in units × Sales
Contribution margin in units = sales price – variable cost
40 – 28 = 12
58,014 = 12 × Q
58,014 / 12 = Q
Q = 4,834.50
Need to increase sales by (4,834.50 – 2,930 =) 1,904.50 units (Answer)

Olori_Odusi Sun, Nov 7, 2021 7:44 AM